OpenStudy (anonymous):

How many Na+ ions are there in 29.25g of NaCl? 3 x 1023 6 x 1023 1.2 x 1026 2.3 x 1023

OpenStudy (anonymous):

I didn't get any of the above. I attached how I solved the question.

OpenStudy (anonymous):

Molar mass of NaCl = 58.5 gm/mol So, no.of moles of NaCl=29.25/58.5=1/2 moles

OpenStudy (anonymous):

now, there are also 1/2 moles of Na+ ion in 1/2 moles of NaCl... So, No. of Na+ ions = 1/2 * 6.23 * 10^23 = 3*10^23

OpenStudy (anonymous):

thanx everyone :)